# appendix,by

### v1

@@ 4.1 Array Modelling and submultiplicativity of burden
Let $\lambda$ be a nonzero limit ordinal and let $A$ be as in the statement of the proposition. We may assume that $E=\emptyset$. We will inductively replace the rows of $A$, making the new rows mutually indiscernible over the remaining ones and keeping the array $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $A$. More formally, we construct recursively on $\alpha<\lambda$ arrays $A^{\alpha}=(a^{\alpha}_{ij})_{i<\lambda,j<\omega}$ such that
@@
    $A^{\alpha}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $A$, and
@@
    $a^{\alpha}_{<\alpha,<\omega}$ is $\mathcal{L}_{\alpha}$-indiscernible over $a^{\alpha}_{\geq\alpha,<\omega}$.
@@
First, let us argue that this is enough to complete the limit step of the induction. Suppose we have $A^{\alpha}$ as above, and put $b_{ij}\coloneqq a^{i+1}_{ij}$ for all $i<\lambda$ and $j<\omega$. To see that $b_{<\lambda,<\omega}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $A$, consider any $\varphi(x_{\bar{\nu}})$ in the $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-type of $A$. Then there is $\alpha<\lambda$ such that $\bar{\nu}\subseteq\alpha\times\omega$. As $A^{\alpha}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $A$, we get that $\vDash\varphi(a^{\alpha}_{\bar{\nu}})$, but $a^{\alpha}_{\bar{\nu}}=b_{\bar{\nu}}$, so $\vDash\varphi(b_{\bar{\nu}})$. Similarly, to see that $(b_{ij})_{i<\lambda,j<\omega}$ is $\mathcal{L}_{\lambda}$-indiscernible, let $\bar{\nu},\bar{\nu}^{\prime}\subseteq\lambda\times\omega$ be finite tuples with $\operatorname{qftp}_{\mathcal{L}_{\lambda}}(\bar{\nu})=\operatorname{qftp}_{\mathcal{L}_{\lambda}}(\bar{\nu}^{\prime})$ and choose $\alpha<\lambda$ such that $\bar{\nu},\bar{\nu}^{\prime}\subseteq\alpha\times\omega$. As $\bar{a}^{\alpha}_{<\alpha}$ is in particular $\mathcal{L}_{\alpha}$-indiscernible over $\emptyset$, we get that $a^{\alpha}_{\bar{\nu}}\equiv a^{\alpha}_{\bar{\nu}^{\prime}}$, i.e. $b_{\bar{\nu}}\equiv b_{\bar{\nu}^{\prime}}$.
@@
We now carry out the construction of the arrays $A^{\alpha}$. If $\alpha<\lambda$ is a limit ordinal, we put $a^{\alpha}_{ij}\coloneqq a^{i+1}_{ij}$ for every $i<\alpha$ and $a^{\alpha}_{ij}=a_{ij}$ for every $i\geq\alpha$. Then (1) and (2) are clearly satisfied by this choice and the inductive assumption, and (3) and (4) are satisfied as well, due to the inductive assumption, the fact that every finite tuple of elements of $\alpha\times\omega$ is contained in $\beta\times\omega$ for some $\beta<\alpha$, and arguments similar to those in the previous paragraph.
@@
For the successor step of the construction, suppose we have constructed $A^{\alpha}$ satisfying conditions (1)(4). Let $)\kappa\coloneqq\beth((2^{\lambda+\lvert A\rvert+\lvert T\rvert}$$^{+})$ and let $\bar{c}=(c_{i})_{i<\kappa}$ be a sequence $\mathrm{EM}_{<}$-based on $\bar{a}_{\alpha}$ over $\bar{a}^{\alpha}_{\neq\alpha}$, which exists by compactness. [^16] Note that we do not need to require this sequence to be indiscernible; this could nonetheless be ensured by invoking 4.4. By Lemma˜4.7, the array $\bar{a}^{\alpha}_{<\alpha}\frown\bar{c}\frown\bar{a}^{\alpha}_{>\alpha}$ (obtained by replacing the $\alpha$-th row of $A^{\alpha}$ with $\bar{c}$) is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $A^{\alpha}$, and hence on $A$ by inductive assumption on $\alpha$. By the main inductive assumption, there is an array $(\tilde{a}^{\alpha}_{ij})_{i<\alpha}$ which is $\mathrm{EM}_{\mathcal{L}_{\alpha}}$-based on $\bar{a}^{\alpha}_{<\alpha}$ over $\bar{a}^{\alpha}_{>\alpha},\bar{c}$ and $\mathcal{L}_{\alpha}$}-indiscernible over $\bar{a}^{\alpha}_{>\alpha},\bar{c}$. By Lemma˜4.7 again, $\tilde{a}^{\alpha}_{<\alpha}\frown\bar{c}\frown\bar{a}^{\alpha}_{>\alpha}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $\bar{a}^{\alpha}_{<\alpha}\frown\bar{c}\frown\bar{a}^{\alpha}_{>\alpha}$, and hence on $A$. As $\bar{a}^{\alpha}_{<\alpha}$ is $\mathcal{L}_{\alpha}$}-indiscernible over $\bar{a}^{\alpha}_{>\alpha}$, we have $\operatorname{EM}_{\mathcal{L}_{\lambda}}(\bar{a}^{\alpha}_{<\alpha}/\bar{a}^{\alpha}_{>\alpha},\bar{c})\supseteq\operatorname{tp}(\bar{a}^{\alpha}_{<\alpha}/\bar{a}^{\alpha}_{>\alpha})$, so $\tilde{a}^{\alpha}_{<\alpha}\bar{a}^{\alpha}_{>\alpha}\equiv\bar{a}^{\alpha}_{<\alpha}\bar{a}^{\alpha}_{>\alpha}$. Let $f$ be an automorphism over $\bar{a}^{\alpha}_{>\alpha}$ sending $\tilde{a}^{\alpha}_{<\alpha}$ to $\bar{a}^{\alpha}_{<\alpha}$. Then $\bar{a}^{\alpha}_{<\alpha}$ is $\mathcal{L}_{\alpha}$}-indiscernible over $f(\bar{c}),\bar{a}^{\alpha}_{>\alpha}$, and $\bar{a}^{\alpha}_{<\alpha}\frown f(\bar{c})\frown\bar{a}^{\alpha}_{>\alpha}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $A$.
@@
Now by ˜4.3 we can choose $\bar{d}=(d_{j})_{j<\omega}$ to be $\{<\}$-based on $f(\bar{c})$ over $\bar{a}^{\alpha}_{<\alpha},\bar{a}^{\alpha}_{>\alpha}$ and indiscernible over $\bar{a}^{\alpha}_{<\alpha},\bar{a}^{\alpha}_{>\alpha}$. By Lemma˜4.7, $\bar{a}^{\alpha}_{<\alpha}\frown\bar{d}\frown\bar{a}^{\alpha}_{>\alpha}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $\bar{a}^{\alpha}_{<\alpha}\frown f(\bar{c})\frown\bar{a}^{\alpha}_{>\alpha}$, and hence on $A$. Note that, since $\bar{a}^{\alpha}_{<\alpha}$ is $\mathcal{L}_{\alpha}$-indiscernible over $f(\bar{c}),\bar{a}^{\alpha}_{>\alpha}$, by Lemma˜4.6 it is also $\mathcal{L}_{\alpha}$-indiscernible over $\bar{d},\bar{a}^{\alpha}_{>\alpha}$. As $\bar{d}$ is indiscernible over $\bar{a}^{\alpha}_{<\alpha},\bar{a}^{\alpha}_{>\alpha}$, it follows from Lemma˜4.8 that $\bar{a}^{\alpha}_{<\alpha}\frown\bar{d}$ is $\mathcal{L}_{\alpha+1}$-indiscernible over $\bar{a}^{\alpha}_{>\alpha}$. Thus putting $a^{\alpha+1}_{ij}\coloneqq a^{\alpha}_{ij}$ for $i\neq\alpha$, and $a^{\alpha+1}_{\alpha,j}\coloneqq d_{j}$ for all $j<\omega$, we get that $A^{\alpha+1}$ satisfies conditions (3) and (4). It also clearly satisfies conditions (1) and (2) by the inductive assumption about $A^{\alpha}$ and the choice of $a^{\alpha+1}_{ij}$. This finishes the recursive construction of $(A^{\alpha})_{\alpha<\lambda}$, and hence the limit step of the inductive proof of the proposition.
@@
We proceed similarly as in the successor step of the above construction. Namely, suppose the statement holds for some $\lambda$ and consider an array $(a_{ij})_{i<\lambda+1,j<\omega}$. We may assume that $E=\emptyset$. By compactness there is a sufficiently long sequence $\bar{c}$ that is $\mathrm{EM}_{<}$-based on $\bar{a}_{\lambda}$ over $\bar{a}_{<\lambda}$. By the inductive hypothesis, there is an $\mathcal{L}_{\lambda}$-indiscernible over $\bar{c}$ array $(b_{ij})_{i<\lambda,j<\omega}$ that is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $\bar{a}_{<\lambda}$ over $\bar{c}$.
Use ˜4.3 to find a $\bar{b}_{<\lambda}$-indiscernible sequence $(b_{\lambda,i})_{i<\omega}$ that is based on $\bar{c}$ over $\bar{b}_{<\lambda}$. By Lemma˜4.6, $\bar{b}_{<\lambda}$ is $\mathcal{L}_{\lambda}$-indiscernible over $\bar{b}_{\lambda}$, hence we may apply Lemma˜4.8 and conclude that $\bar{b}_{<\lambda+1}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda+1}}$-indiscernible.
We conclude by observing that, by repeated applications of Lemma˜4.7, $\bar{b}_{<\lambda+1}$ is $\mathrm{EM}_{\mathcal{L}_{\lambda+1}}$-based on $\bar{b}_{<\lambda}\frown\bar{c}$, in turn $\mathrm{EM}_{\mathcal{L}_{\lambda+1}}$-based on $\bar{a}_{<\lambda}\frown\bar{c}$, itself $\mathrm{EM}_{\mathcal{L}_{\lambda+1}}$-based on $\bar{a}_{<\lambda+1}$. ∎
@@
Without loss of generality $E=\emptyset$. Let $)\lambda\coloneqq\beth((2^{\lvert T\rvert+\lvert a_{0,0}\rvert}$$^{+})$. By compactness we can extend $(a_{ij})_{i,j<\omega}$ to an array $(a_{ij})_{i<\lambda,j<\omega}$ which is $\mathrm{EM}_{\mathcal{L}_{\mathrm{ar}}}$- based on $(a_{ij})_{i,j<\omega}$. By Proposition˜4.9, there is an $\mathcal{L}_{\lambda}$-indiscernible array $(c_{ij})_{i<\lambda,j<\omega}$ which is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$-based on $(a_{ij})_{i<\lambda,j<\omega}$. In particular, $(c_{ij})_{i<\lambda,j<\omega}$ is $\mathrm{EM}_{\mathcal{L}_{\mathrm{ar}}}$-based on $(a_{ij})_{i<\lambda,j<\omega}$. By ˜4.3 there is an indiscernible sequence $(\bar{b}_{i})_{i<\omega}$ based on the sequence $\bar{c}_{<\lambda}$. Then, by ˜4.5, $(b_{ij})_{i,j<\omega}$ is an $\mathcal{L}_{\mathrm{ar}}$-indiscernible array $\mathrm{EM}_{\mathcal{L}_{\mathrm{ar}}}$-based on $(a_{ij})_{i,j<\omega}$. ∎
In [Kam24], the thickness assumption was used in proving Array Modelling, but not in deriving its consequences. In other words, having proved Array Modelling without assuming thickness, we may derive, by the same proofs as in [Kam24, Lemma 6.2 and Theorem 1.4], the following corollaries.
@@
    for every $\alpha<\kappa$ there are $k_{\alpha}<\omega$ and $\psi_{\alpha}(y_{0},\ldots,y_{k_{\alpha}-1})$ such that
    $\vDash\neg\left(\exists x,y_{0},\ldots,y_{k_{\alpha}-1}\;\left(\psi_{\alpha}(y_{0},\ldots,y_{k_{\alpha}-1})\wedge\bigwedge_{i<k_{\alpha}}\varphi_{\alpha}(x,y_{i})\right)\right)$
@@
Following [Che14], we will say an array $(a_{ij})_{i<\alpha,j<\beta}$ is almost mutually indiscernible over $E$ iff for every $i_{0}<\alpha$ we have that $(a_{i_{0},j})_{j<\beta}$ is an indiscernible sequence over $a_{<i_{0},<\beta}a_{>i_{0},0}E$.
In the next lemma we correct [Che14, Lemma 2.3], which we will see in ˜4.20 not to hold as stated in [Che14].
@@
We may assume $E=\emptyset$. Let $b^{\prime\prime}\vDash p_{\infty}(x)$ and use ˜4.4 to obtain some $\bar{a}^{\prime\prime}$ indiscernible over $b^{\prime\prime}$ and $\mathrm{EM}_{<}$-based on $\bar{a}$ over $b^{\prime\prime}$.
Let $f$ be an automorphism sending $\bar{a}^{\prime\prime}$ to $\bar{a}$ and put $b^{\prime}\coloneqq f(b^{\prime\prime})$. Note that $b^{\prime}a_{0}\equiv b^{\prime\prime}a^{\prime\prime}_{0}$. Moreover, $b^{\prime\prime}a^{\prime\prime}_{0}\equiv ba_{0}$, because for every $i$ we have $b^{\prime\prime}a_{i}\equiv ba_{0}$ by choice of $b^{\prime\prime}$. Therefore, there is an automorphism $g$ sending $b^{\prime}a_{0}$ to $ba_{0}$. Put $\bar{a}^{\prime}\coloneqq g(\bar{a})$. Then, clearly, $\bar{a}^{\prime}\equiv_{a_{0}}\bar{a}$, and, as $\bar{a}^{\prime\prime}$ is indiscernible over $b^{\prime\prime}$, we get that $\bar{a}^{\prime}=g(f(\bar{a}^{\prime\prime}))$ is indiscernible over $g(f(b^{\prime\prime}))=g(b^{\prime})=b$. ∎
@@
    $\bar{a}_{\alpha}^{\prime}$ is indiscernible over $b_{2}\bar{a}_{<\alpha}^{\prime}\bar{a}_{\geq(\alpha+1,0)}$.
@@
    $\bar{a}_{\alpha}^{\prime}$ has the same type as $\bar{a}_{(\alpha,\beta_{\alpha})}$ over $a_{(\alpha,\beta_{\alpha}),0}\bar{a}_{<\alpha}^{\prime}\bar{a}_{\geq(\alpha+1,0)}$.
@@
We now provide a counterexample to [Che14, Lemma 2.3], and also to its relaxation to “almost mutually indiscernible”. The lemma is used in that paper only with $\kappa=1$ in which case it is true (it is Lemma˜4.15 above, specialised to first-order logic), hence this does not affect correctness of any other results in there.
@@
Let $\mathcal{L}$ be the language with three sorts $\mathrm{A},\mathrm{B},\mathrm{C}$ and a single function symbol $f\colon\mathrm{A}\times\mathrm{B}\to\mathrm{C}$. One checks amalgamation of finite $\mathcal{L}$-structures easily, obtaining a Fraïssé class. Let $T$ be the theory of its Fraïssé limit (see [HM23, Definition 1.2(viii)] for an axiomatisation).
Let $\bar{c}=(c_{i})_{i<\omega}$ be a sequence of pairwise distinct elements of sort $\mathrm{C}$, and let $a_{0}$ be an arbitrary element of sort $\mathrm{A}$. Consider the $2$-row array with first row $\bar{c}$ and second row $\bar{a}$ constantly equal to $a_{0}$. Let $C=\emptyset$. This is a counterexample to [Che14, Lemma 2.3].
@@
Choose any $b\in\mathrm{B}$ such that $f(a_{0},b)=c_{0}$. By quantifier elimination and genericity of $f$, the partial type $p^{\infty}$ in [Che14, Lemma 2.3] is consistent (in fact, it defines the whole sort $\mathrm{B}$).

### v2

@@ 4.1 Array Modelling and submultiplicativity of burden
Let $\lambda$ be a nonzero limit ordinal and let $A$ be as in the statement of the proposition. We may assume that $E=\emptyset$. We will inductively replace the rows of $A$, making the new rows mutually indiscernible over the remaining ones and keeping the array $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $A$. More formally, we construct recursively on $\alpha<\lambda$ arrays $A^{\alpha}=(a^{\alpha}_{ij})_{i<\lambda,j<\omega}$ such that
@@
    $A^{\alpha}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $A$, and
@@
    $a^{\alpha}_{<\alpha,<\omega}$ is $\mathfrak{L}_{\alpha}$-indiscernible over $a^{\alpha}_{\geq\alpha,<\omega}$.
@@
First, let us argue that this is enough to complete the limit step of the induction. Suppose we have $A^{\alpha}$ as above, and put $b_{ij}\coloneqq a^{i+1}_{ij}$ for all $i<\lambda$ and $j<\omega$. To see that $b_{<\lambda,<\omega}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $A$, consider any $\varphi(x_{\bar{\nu}})$ in the $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-type of $A$. Then there is $\alpha<\lambda$ such that $\bar{\nu}\subseteq\alpha\times\omega$. As $A^{\alpha}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $A$, we get that $\vDash\varphi(a^{\alpha}_{\bar{\nu}})$, but $a^{\alpha}_{\bar{\nu}}=b_{\bar{\nu}}$, so $\vDash\varphi(b_{\bar{\nu}})$. Similarly, to see that $(b_{ij})_{i<\lambda,j<\omega}$ is $\mathfrak{L}_{\lambda}$-indiscernible, let $\bar{\nu},\bar{\nu}^{\prime}\subseteq\lambda\times\omega$ be finite tuples with $\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\bar{\nu})=\operatorname{qftp}_{\mathfrak{L}_{\lambda}}(\bar{\nu}^{\prime})$ and choose $\alpha<\lambda$ such that $\bar{\nu},\bar{\nu}^{\prime}\subseteq\alpha\times\omega$. As $\bar{a}^{\alpha}_{<\alpha}$ is in particular $\mathfrak{L}_{\alpha}$-indiscernible over $\emptyset$, we get that $a^{\alpha}_{\bar{\nu}}\equiv a^{\alpha}_{\bar{\nu}^{\prime}}$, i.e. $b_{\bar{\nu}}\equiv b_{\bar{\nu}^{\prime}}$.
@@
We now carry out the construction of the arrays $A^{\alpha}$. If $\alpha<\lambda$ is a limit ordinal, we put $a^{\alpha}_{ij}\coloneqq a^{i+1}_{ij}$ for $i<\alpha$ and $a^{\alpha}_{ij}=a_{ij}$ for $i\geq\alpha$. Then (1) and (2) are clearly satisfied by this choice and the inductive assumption, and (3) and (4) are satisfied as well, due to the inductive assumption, the fact that every finite tuple of elements of $\alpha\times\omega$ is contained in $\beta\times\omega$ for some $\beta<\alpha$, and arguments similar to those in the previous paragraph.
@@
For the successor step of the construction, suppose we have constructed $A^{\alpha}$ satisfying conditions (1)(4). Let $)\kappa\coloneqq\beth((2^{\lambda+\lvert A\rvert+\lvert T\rvert}$$^{+})$ and let $\bar{c}=(c_{i})_{i<\kappa}$ be a sequence $\mathrm{EM}_{<}$-based on $\bar{a}_{\alpha}$ over $\bar{a}^{\alpha}_{\neq\alpha}$, which exists by compactness. [^16] Note that we do not need to require this sequence to be indiscernible; this could nonetheless be ensured by invoking 4.4. By Lemma˜4.7, the array $\bar{a}^{\alpha}_{<\alpha}\frown\bar{c}\frown\bar{a}^{\alpha}_{>\alpha}$ (obtained by replacing the $\alpha$-th row of $A^{\alpha}$ with $\bar{c}$) is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $A^{\alpha}$, and hence on $A$ by inductive assumption on $\alpha$. By the main inductive assumption, there is an array $(\hat{a}^{\alpha}_{ij})_{i<\alpha}$ which is $\mathrm{EM}_{\mathfrak{L}_{\alpha}}$-based on $\bar{a}^{\alpha}_{<\alpha}$ over $\bar{a}^{\alpha}_{>\alpha},\bar{c}$ and $\mathfrak{L}_{\alpha}$}-indiscernible over $\bar{a}^{\alpha}_{>\alpha},\bar{c}$. By Lemma˜4.7 again, $\hat{a}^{\alpha}_{<\alpha}\frown\bar{c}\frown\bar{a}^{\alpha}_{>\alpha}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $\bar{a}^{\alpha}_{<\alpha}\frown\bar{c}\frown\bar{a}^{\alpha}_{>\alpha}$, and hence on $A$. As $\bar{a}^{\alpha}_{<\alpha}$ is $\mathfrak{L}_{\alpha}$}-indiscernible over $\bar{a}^{\alpha}_{>\alpha}$, we have $\operatorname{EM}_{\mathfrak{L}_{\lambda}}(\bar{a}^{\alpha}_{<\alpha}/\bar{a}^{\alpha}_{>\alpha},\bar{c})\supseteq\operatorname{tp}(\bar{a}^{\alpha}_{<\alpha}/\bar{a}^{\alpha}_{>\alpha})$, so $\hat{a}^{\alpha}_{<\alpha}\bar{a}^{\alpha}_{>\alpha}\equiv\bar{a}^{\alpha}_{<\alpha}\bar{a}^{\alpha}_{>\alpha}$. Let $f$ be an automorphism over $\bar{a}^{\alpha}_{>\alpha}$ sending $\hat{a}^{\alpha}_{<\alpha}$ to $\bar{a}^{\alpha}_{<\alpha}$. Then $\bar{a}^{\alpha}_{<\alpha}$ is $\mathfrak{L}_{\alpha}$}-indiscernible over $f(\bar{c}),\bar{a}^{\alpha}_{>\alpha}$, and $\bar{a}^{\alpha}_{<\alpha}\frown f(\bar{c})\frown\bar{a}^{\alpha}_{>\alpha}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $A$.
@@
Now by ˜4.3 we can choose $\bar{d}=(d_{j})_{j<\omega}$ to be $\{<\}$-based on $f(\bar{c})$ over $\bar{a}^{\alpha}_{<\alpha},\bar{a}^{\alpha}_{>\alpha}$ and indiscernible over $\bar{a}^{\alpha}_{<\alpha},\bar{a}^{\alpha}_{>\alpha}$. By Lemma˜4.7, $\bar{a}^{\alpha}_{<\alpha}\frown\bar{d}\frown\bar{a}^{\alpha}_{>\alpha}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $\bar{a}^{\alpha}_{<\alpha}\frown f(\bar{c})\frown\bar{a}^{\alpha}_{>\alpha}$, and hence on $A$. Note that, since $\bar{a}^{\alpha}_{<\alpha}$ is $\mathfrak{L}_{\alpha}$-indiscernible over $f(\bar{c}),\bar{a}^{\alpha}_{>\alpha}$, by Lemma˜4.6 it is also $\mathfrak{L}_{\alpha}$-indiscernible over $\bar{d},\bar{a}^{\alpha}_{>\alpha}$. As $\bar{d}$ is indiscernible over $\bar{a}^{\alpha}_{<\alpha},\bar{a}^{\alpha}_{>\alpha}$, it follows from Lemma˜4.8 that $\bar{a}^{\alpha}_{<\alpha}\frown\bar{d}$ is $\mathfrak{L}_{\alpha+1}$-indiscernible over $\bar{a}^{\alpha}_{>\alpha}$. Thus putting $a^{\alpha+1}_{ij}\coloneqq a^{\alpha}_{ij}$ for $i\neq\alpha$, and $a^{\alpha+1}_{\alpha,j}\coloneqq d_{j}$ for all $j<\omega$, we get that $A^{\alpha+1}$ satisfies conditions (3) and (4). It also clearly satisfies conditions (1) and (2) by the inductive assumption about $A^{\alpha}$ and the choice of $a^{\alpha+1}_{ij}$. This finishes the recursive construction of $(A^{\alpha})_{\alpha<\lambda}$, and hence the limit step of the inductive proof of the proposition.
@@
We proceed similarly as in the successor step of the above construction. Namely, suppose the statement holds for some $\lambda$ and consider an array $(a_{ij})_{i<\lambda+1,j<\omega}$. We may assume that $E=\emptyset$. By compactness there is a sufficiently long sequence $\bar{c}$ that is $\mathrm{EM}_{<}$-based on $\bar{a}_{\lambda}$ over $\bar{a}_{<\lambda}$. By the inductive hypothesis, there is an $\mathfrak{L}_{\lambda}$-indiscernible over $\bar{c}$ array $(b_{ij})_{i<\lambda,j<\omega}$ that is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $\bar{a}_{<\lambda}$ over $\bar{c}$.
Use ˜4.3 to find a $\bar{b}_{<\lambda}$-indiscernible sequence $(b_{\lambda,i})_{i<\omega}$ that is based on $\bar{c}$ over $\bar{b}_{<\lambda}$. By Lemma˜4.6, $\bar{b}_{<\lambda}$ is $\mathfrak{L}_{\lambda}$-indiscernible over $\bar{b}_{\lambda}$, hence we may apply Lemma˜4.8 and conclude that $\bar{b}_{<\lambda+1}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda+1}}$-indiscernible.
We conclude by observing that, by repeated applications of Lemma˜4.7, $\bar{b}_{<\lambda+1}$ is $\mathrm{EM}_{\mathfrak{L}_{\lambda+1}}$-based on $\bar{b}_{<\lambda}\frown\bar{c}$, in turn $\mathrm{EM}_{\mathfrak{L}_{\lambda+1}}$-based on $\bar{a}_{<\lambda}\frown\bar{c}$, itself $\mathrm{EM}_{\mathfrak{L}_{\lambda+1}}$-based on $\bar{a}_{<\lambda+1}$. ∎
@@
Without loss of generality $E=\emptyset$. Let $)\lambda\coloneqq\beth((2^{\lvert T\rvert+\lvert a_{0,0}\rvert}$$^{+})$. By compactness we can extend $(a_{ij})_{i,j<\omega}$ to an array $(a_{ij})_{i<\lambda,j<\omega}$ which is $\mathrm{EM}_{\mathfrak{L}_{\mathrm{ar}}}$-based on $(a_{ij})_{i,j<\omega}$. By Proposition˜4.9, there is an $\mathfrak{L}_{\lambda}$-indiscernible array $(c_{ij})_{i<\lambda,j<\omega}$ which is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$-based on $(a_{ij})_{i<\lambda,j<\omega}$. In particular, by ˜4.2(1), $(c_{ij})_{i<\lambda,j<\omega}$ is $\mathrm{EM}_{\mathfrak{L}_{\mathrm{ar}}}$-based on $(a_{ij})_{i<\lambda,j<\omega}$. By ˜4.3 there is an indiscernible sequence $(\bar{b}_{i})_{i<\omega}$ based on the sequence $\bar{c}_{<\lambda}$. Then, by ˜4.5, $(b_{ij})_{i,j<\omega}$ is an $\mathfrak{L}_{\mathrm{ar}}$-indiscernible array $\mathrm{EM}_{\mathfrak{L}_{\mathrm{ar}}}$-based on $(a_{ij})_{i,j<\omega}$. ∎
In [Ka], the thickness assumption was used in proving Array Modelling, but not in deriving its consequences. In other words, having proved Array Modelling without assuming thickness, we may derive, by the same proofs as in [Ka, Lemma 6.2 and Theorem 1.4], the following corollaries.
@@
    for every $\alpha<\kappa$ there are $k_{\alpha}<\omega$ and $\psi_{\alpha}(y_{\alpha,0},\ldots,y_{\alpha,{k_{\alpha}-1}})$ such that
    $\vDash\neg\left(\exists x,y_{\alpha,0},\ldots,y_{\alpha,{k_{\alpha}-1}}\;\left(\psi_{\alpha}(y_{\alpha,0},\ldots,y_{\alpha,{k_{\alpha}-1}})\wedge\bigwedge_{i<k_{\alpha}}\varphi_{\alpha}(x,y_{\alpha,i})\right)\right)$
@@
Following [Chernikov], we will say that an array $(a_{ij})_{i<\alpha,j<\beta}$ is almost mutually indiscernible over $E$ iff for every $i_{0}<\alpha$ we have that $(a_{i_{0},j})_{j<\beta}$ is an indiscernible sequence over $a_{<i_{0},<\beta}a_{>i_{0},0}E$.
In the next lemma we correct [Chernikov, Lemma 2.3], which we will see in ˜4.20 not to hold as stated in [Chernikov].
@@
We may assume $E=\emptyset$. Let $b^{\prime}\vDash p_{\infty}(x)$ and use ˜4.4 to obtain some $\bar{a}^{\prime\prime}$ indiscernible over $b^{\prime}$ and $\mathrm{EM}_{<}$-based on $\bar{a}$ over $b^{\prime}$.
Note that $b^{\prime}a^{\prime\prime}_{0}\equiv ba_{0}$, because for every $i$ we have $b^{\prime}a_{i}\equiv ba_{0}$ by choice of $b^{\prime}$. Therefore, there is an automorphism $f$ sending $b^{\prime}a_{0}^{\prime\prime}$ to $ba_{0}$. Put $\bar{a}^{\prime}\coloneqq f(\bar{a}^{\prime\prime})$. Clearly $\bar{a}^{\prime}\equiv_{a_{0}}\bar{a}$ and, as $\bar{a}^{\prime\prime}$ is indiscernible over $b^{\prime}$, we get that $\bar{a}^{\prime}=f(\bar{a}^{\prime\prime})$ is indiscernible over $b=f(b^{\prime})$. ∎
@@
    $\bar{a}_{\alpha}^{\prime}$ is indiscernible over $b_{2}\bar{a}_{<\alpha}^{\prime}\bar{a}_{\geq(\alpha+1,0)}$,
@@
    $\bar{a}_{\alpha}^{\prime}$ has the same type as $\bar{a}_{(\alpha,\beta_{\alpha})}$ over $a_{(\alpha,\beta_{\alpha}),0}\bar{a}_{<\alpha}^{\prime}\bar{a}_{\geq(\alpha+1,0)}$, and
@@
We now provide a counterexample to [Chernikov, Lemma 2.3], and also to its relaxation to “almost mutually indiscernible”. The lemma is used in that paper only with $\kappa=1$ in which case it is true (it is Lemma˜4.15 above, specialised to first-order logic), hence this does not affect correctness of any other results in there.
@@
Let $\mathcal{L}$ be the language with three sorts $\mathrm{A},\mathrm{B},\mathrm{C}$ and a single function symbol $f\colon\mathrm{A}\times\mathrm{B}\to\mathrm{C}$. One checks amalgamation of finite $\mathcal{L}$-structures easily, obtaining a Fraïssé class. Let $T$ be the theory of its Fraïssé limit (see [HM, Definition 1.2(viii)] for an axiomatisation).
Let $\bar{c}=(c_{i})_{i<\omega}$ be a sequence of pairwise distinct elements of sort $\mathrm{C}$, and let $a_{0}$ be an arbitrary element of sort $\mathrm{A}$. Consider the $2$-row array with first row $\bar{c}$ and second row $\bar{a}$ constantly equal to $a_{0}$. Let $C=\emptyset$. This is a counterexample to [Chernikov, Lemma 2.3].
@@
Choose any $b\in\mathrm{B}$ such that $f(a_{0},b)=c_{0}$. By quantifier elimination and genericity of $f$, the partial type $p^{\infty}$ in [Chernikov, Lemma 2.3] is consistent (in fact, it defines the whole sort $\mathrm{B}$).